OpenStudy (anonymous):

Write an equation of a line with the given slope and y-intercept. m = –5, b = –3

OpenStudy (anonymous):

general equation for straight line: y=m*x+c Because the line intercept y at -3, it pass through (0,-3) -3=0+c or c=-3 eq: y=-5x-3

OpenStudy (anonymous):

^_^ thanks